OpenStudy (anonymous):

Use Gauss's approach to find the following sums. (Do not use formulas.) a. 1+2+3+4+...+99 b. 1+3+5+7+...+103

OpenStudy (anonymous):

i don't know Gauss approach. 1+2+3+............97+98+99 =(1+99)+(2+98)+(3+97)+(4+96)+.....+(48+52)+(49+51)+50 =100*49+50=4950

OpenStudy (anonymous):

Okay thank you so much! I am not very good at Math and did not understand.

OpenStudy (anonymous):

I need the sum of a, and b

OpenStudy (anonymous):

i have not used formula ,only arranged the terms.

OpenStudy (anonymous):

first term goes on increasing second goes on decreasing.

OpenStudy (anonymous):

b. 1+3+5+...+99+101+103 =1+2+3+...+99-(2+4+6+...+98)+101+103 =4950-2(1+2+3+...+49)+204 =4950-2{(1+49)+(2+48)+(3+47)+...+(23+27)+(24+26)+25}+204 =4950-2(50*24+25)+204 =4950-2*1225+204 =4950-2450+204 =2500+204 =2704